The City of Wolfe has issued its financial statements for Year 4 (assume that the city uses a calendar year). The citys general fund is made up of two functions: (1) education and (2) parks. The city also utilizes capital projects funds for ongoing construction and an enterprise fund to account for an art museum. It also has one discretely presented component unit. The government-wide financial statements indicated the following Year 4 totals: Education had net expenses of $710,000. Parks had net expenses of $130,000. Art museum had net revenues of $80,000 General revenues were $900,000; the overall increase in net position was $140,000. The fund financial statements issued for Year 4 indicated the following The general fund had an increase of $30,000 in its fund balance The capital projects fund had an increase of $40,000 in its fund balance The enterprise fund had an increase of $60,000 in its net position. Officials for Wolfe define available as current financial resources to be paid or collected within 60 days. On the first day of Year 4, the city receives a painting as a gift that qualifies as a work of art. It has a 30-year life, is worth $15,000, and is being displayed at one of the local parks. The accountant accidentally capitalized and depreciated it although officials had wanted to use the allowed alternative Respond to the following questions a. According to the information provided above, the general fund reported a $30,000 increase in its fund balance. If city officials had used proper alternatives in this reporting, what would have been the correct change in the fund balance for the general fund for the year? alternatives in this reporting, what was the correct net expense for parks for the year? overall change in net position for Year 4 on government-wide financial statements, assuming that officials still preferred the allowed b. According to the information provided above, the parks reported net expenses of $130,000. If city officials had used proper c. Assume the same information except that the art was given to the art museum but not recorded at all. What should have been the alternative? a. The correct increase in fund balance is b. Net expense c. The correct increase in net asset is $30,000
Answer #1

Answer: A $30000

Fund financial statements are not affected by gift. Only government wide statements are affected by the asset and depreciation. The increase of $30000 is appropriate and should be changed. Therefore, the correct increase in fund balance = $30000

Answer: B $ 144,500

Increase in net position = Revenue – depreciation recognized for the first year

= 15000-(15000/30) = 14500

Net expense = 130000+14500 = 144500

Answer: C $ 140,000

If expense is recorded along with revenue of $15000, then there would be no change in the net position it will be remain same that is $140000

Therefore,

a.

The correct increase in fund balance

$30000

b.

Net expense

$ 144,500

c.

The correct increase in net asset is

$ 140,000